OpenStudy (anonymous):

Hello. I need to know how to solve: find the vertex of the equation: y=-3x^2-4x+9.plueze help!

OpenStudy (saifoo.khan):

\[Vertex = (-\frac23, \frac{31}{2})\]

OpenStudy (anonymous):

thanks. would you show the work? I need to know step by step. thanks.

OpenStudy (precal):

there is a formula for the vertex.

OpenStudy (saifoo.khan):

Shortcut: -b/2a. \[-\frac{-4}{2(-3)}\to - \frac{-4}{-6} \to -\frac12\]

OpenStudy (saifoo.khan):

Now insert this value in the equation and find out the y-cordinate.

OpenStudy (anonymous):

we just over this in class last night. so pardon me for saying...where in the equation do i insert it?

OpenStudy (precal):

in the original

OpenStudy (anonymous):

so I subsitute the x?

OpenStudy (precal):

yes

OpenStudy (anonymous):

okay thanks

OpenStudy (saifoo.khan):

Have more problems?

OpenStudy (anonymous):

lots.

OpenStudy (saifoo.khan):

Post em!

OpenStudy (anonymous):

actually 4

OpenStudy (saifoo.khan):

Wow, we will take ages to do those.

OpenStudy (anonymous):

lol

OpenStudy (saifoo.khan):

;)

OpenStudy (anonymous):

quadratic equations with solutions of \[{ 3/4 \-1/4}\] is defined as: write the equation

OpenStudy (saifoo.khan):

\[\frac34 , -\frac14\]That?

OpenStudy (anonymous):

yes

OpenStudy (precal):

(x-3/4)(x+1/4)

OpenStudy (precal):

multiply it out

OpenStudy (anonymous):

okay